OpenStudy (anonymous):

Verify the given linear approximation at a = 0. Then determine the values of x for which the linear approximation is accurate to within 0.1. (Enter your answer using interval notation. Round your answers to three decimal places.) ln(1 + x) ≈ x

OpenStudy (anonymous):

I have no idea how to do this one. I think I need to use this formula: L(x)=f(a)+f'(a) (x-a)

OpenStudy (anonymous):

it is the same as a problem you probably did in the second or third week of class find the equation of the line tangent to the curve of \(y=f(x)\) at \((a,f(a))\)

OpenStudy (anonymous):

so...would it be find the derivative of ln(1+x)=x??

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

they are giving you the answer

OpenStudy (anonymous):

they are telling you that the equation of the line tangent to \(y=\ln(1+x)\) at \((0,1)\) is \(y=x\) and asking you to "check" that it is right

OpenStudy (anonymous):

okay that was wrong, it is at \((0,0)\)

OpenStudy (anonymous):

so, should I just plug in numbers 0.1 greater and smaller into the formula? I really have no idea what to do

OpenStudy (anonymous):

in any case the check is to take the derivative of \(\ln(x+1)\) get \(\frac{1}{x+1}\) then replace \(x\) by \(0\) to find the slope, you get that the slope is 1

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

and since the line goes through the origin, the equation of the tangent line at \((0,0)\) is \(y=x\)

OpenStudy (anonymous):

wait! so would the answer be \[x \in \mathbb{R} \] ??

OpenStudy (anonymous):

no the linear approximation is only good locally, not globally

OpenStudy (anonymous):

I'm sorry, I don't know what to do next, I understand everything else you've done but I dont know how to get what they are asking for

OpenStudy (anonymous):

okay first here is a picture, then we can see how for what values of x we get a small error http://www.wolframalpha.com/input/?i=ln%28x%2B1%29%2C+x+++domain+-0.2..0.2

OpenStudy (anonymous):

without using power series i am not sure exactly how to answer the last question you want to make \(|\ln(x+1)-x|<0.1\) or \(\ln(x+1)-.1<x<\ln(x+1)+.1\)

OpenStudy (anonymous):

so i guess you can just check with numbers and see what works

OpenStudy (anonymous):

any number?? or shouls I use small decimals?

OpenStudy (anonymous):

small ones i don't think this is that easy

OpenStudy (anonymous):

ok, thanks!

OpenStudy (anonymous):

yw, sorry i couldn't be more help